OpenStudy (anonymous):

If P(B) = 0.5, P(A or B) = 0.3, and P(A and B) = 0.6. P(A) =

OpenStudy (anonymous):

no hold on lol

OpenStudy (anonymous):

k

myininaya (myininaya):

\[P(A)+P(B)-P(A \cap B)=P(A \cup B)\]

myininaya (myininaya):

plug in and solve for P(A)

myininaya (myininaya):

\[P(A)+0.5-0.6=0.3\]

OpenStudy (anonymous):

what branch of math is this?

myininaya (myininaya):

probability

OpenStudy (anonymous):

Oh. I was thinking it was finding a function where that is probably cuped.